1 1/2 devid by 1 3/8
Answer by hyperasad(6)

First of all convert proper fractions to improper fractions.
1 1/2 = 3/2 and 1 3/8 = 11/8
now perform division function
(3/2)/ (11/8)
It becomes 3/2 x 8/11
= 3/1 x 4/11
=12/11
answer is in improper fraction. convert it into proper fraction
=1 1/11